### 10 Apr 2026 v2

# Implementation of a shooting technique for quantum unitary control on spin qudits

Abstract
High-fidelity control of quantum systems is essential for scalable quantum technologies. We introduce a shooting-based method which yields smooth control pulses designed to implement gates on discrete quantum systems, and demonstrate its performances through numerical simulations on systems inspired from single molecule magnets. Our method efficiently decomposes quantum gates into electromagnetic pulses, and determines control pulses which are faster than GRAPE, all the more as the system’s dimension increases.
I Introduction
One of the key challenges in advancing quantum simulations and computations is scaling computational power without compromising robustness against noise. Quantum optimal control provides a framework for steering quantum systems toward desired target states with high fidelity in the presence of physical constraints and noise [13, 25]. By shaping external control fields, these methods enable precise manipulation of quantum dynamics. In this work, we provide the description and a numerical benchmark for a quantum optimal control algorithm based on a gradient descent and a shooting technique [6, 31], on a concrete physical system: molecular spin qudits. By combining one or more transition metals or lanthanide ions with total spin greater than 1/2, it is possible to generate multiple quantum states. In appropriately designed molecules, the energies of these states are non-equidistant, allowing selective transitions to be addressed by microwave pulses [10]. In this work, we focus on arrangements of terbium ions between layers (called “deckers” in the literature) of phthalocyanine (Pc) ligands [23]. However, this study may be extended to any process involving an addressable Hilbert space of finite dimension, for instance nitrogen-vacancy in diamond centers [20], NMR systems, trapped ions, Rydberg atoms.
To implement quantum gates on the aforementioned systems, we have to unravel the gates into a sequence of electromagnetic signals sent to the system. Multiple algorithms have been designed to address this issue and most notably the widely used Gradient Ascent Pulse Engineering (GRAPE). This paper showcases a shooting-based method by comparing it to GRAPE on two physically relevant examples, the double and triple decker molecules, $\text{TbPc}_{2}$ and $\text{Tb}_{2}\text{Pc}_{3}$. Our numerical results demonstrate that this shooting method, achieves the desired fidelity level and provides faster execution times, up to 37% for the triple-decker molecule in the case of a quantum Fourier transform for instance. We denote by execution time the actual time it takes to run the gate on a quantum platform and not the (classical) time needed to find the pulses. We chose to numerically compare the results of these two methods, since both can be used without assuming specialized knowledge on the system and implement gradient-based optimisations at their cores. Other methods exist, for instance derived from the work of Krotov [26, 15], which has been compared to GRAPE and other methods in [27].
The section II will take advantage of a presentation of the usual methods for pulse implementation to introduce the framework we will work with in section II.1 and present the developped method in section II.2. We will apply it in section III to the physical system of the $\text{TbPc}_{2}$ and $\text{Tb}_{2}\text{Pc}_{3}$ molecules presented in section III.1 and show the results in section III.2.
II Presentation of the method
II.1 Quantum optimal control of pulses
In several platforms, quantum optimal control of pulses to shape quantum gates are made out of Arbitrary Waweform Generator (AWG), that can accurately control a frequency $\omega$ and a phase $\theta$ for each transition $n$, and each pulse timestep $i$ having duration $\tau$, where $A$ is the amplitude of the pulse. Using $\Theta$ as the Heaviside function, the pulse function $S(t)$ is represented here by:
$S(t)=\sum_{n,i}\Theta(t-t_{n,i})\Theta(t-t_{n,i}-\tau_{n,i})A_{n,i}\cos(\omega_{n}t-\theta_{n,i}).$ (1)
We consider a simple system of two levels, with a drift Hamiltonian $H_{0}$ evolving according to a precession $\omega_{\text{prec}}$ around the z-axis. We can represent the AWG effect by adding a monochromatic electromagnetic field (in a direction orthogonal to the precession, the x-axis for instance) interacting with the magnetic moment of the system [14], which can be described by the following Hamiltonian:
$H$ $=H_{0}+g\overrightarrow{\mu_{N}}\overrightarrow{E}$ (2)
$=\frac{\omega_{\text{prec}}}{2}\sigma_{z}+\Omega_{\text{Rabi}}\cos(\omega t-\theta)\sigma_{x}.$
By choosing a rotating frame with frequency $\omega_{\text{rot}}$, this Hamiltonian can be rewritten as
$H^{\prime}$ $=\frac{(\omega_{\text{prec}}-\omega_{\text{rot}})}{2}\sigma_{z}+\Omega_{\text{Rabi}}\left[(\cos((\omega+\omega_{\text{rot}})t-\theta)+\cos((\omega-\omega_{\text{rot}})t-\theta))\sigma_{x}\right$. (3)
$+\left.(\sin((\omega+\omega_{\text{rot}})t-\theta)-\sin((\omega-\omega_{\text{rot}})t-\theta))\sigma_{y}\right].$
Usually three simplifications are made, choosing $\omega_{\text{rot}}=\omega_{\text{prec}}$ to remove the drift term which places us in the interaction picture, choosing $\omega=\omega_{\text{rot}}$, which places us in the rotating frame, and applying the Rotating Wave Approcimation (RWA) to neglect fast oscillating terms, one ends up with this expression of the Hamiltonian:
$H=\Omega_{\text{Rabi}}\left[\cos(\theta)\sigma_{x}-\sin(\theta)\sigma_{y}\right].$ (4)
In the following of the paper, we place ourselves in this situation where we can apply eq. 4 between two energy levels of the Hilbert spaces we are working with as long as they are considered “connected”, that is as long as such an interaction is experimentally allowed. We can control both $\Omega_{\text{Rabi}}$ and $\theta$ through an AWG.
II.1.1 An analytical method: the Givens Rotation Decomposition (GRD)
The Givens Rotation Decomposition [12] (GRD) is a method that systematically eliminates subdiagonal elements by applying successive rotation matrices operating on some two-dimensional subspace to decompose any matrix into a “QR” decompostion [11] with $Q$ unitary and $R$ upper triangular. For a unitary matrix this amounts to decomposing the matrix into $Q$ unitary and $R$ diagonal. A Givens Rotation (GR), defined by two real parameters $\theta$ and $\phi$ as
$\text{GR}(\phi,\theta)=e^{-i(\cos(\theta)\sigma_{x}+\sin(\theta)\sigma_{y})\phi}$ $=\begin{bmatrix}\cos(\phi)&-ie^{-i\theta}\sin(\phi)\\ -ie^{i\theta}\sin(\phi)&\cos(\phi)\end{bmatrix}$, (5)
is embedded in a larger space to be used to zero out an element of a column (See the first step in eq. 11), and succesive rotations allow to zero out all subdiagonal elements for a given column (The second step in eq. 11). Then we can do it on the next column, without removing the zeros on the previous columns since the GR applied on zeros does not modify them.
$U$ $\xrightarrow[]{}\left(\begin{array}[]{cccc}u_{1,1}&u_{1,2}&\cdots&u_{1,N}\\ \vdots&\vdots&\ddots&\vdots\\ u_{N-2,1}&u_{N-2,2}&\cdots&u_{N-2,N}\\ \tilde{u}_{N-1,1}&\tilde{u}_{N-1,2}&\cdots&\tilde{u}_{N-1,N}\\ 0&\tilde{u}_{N,2}&\cdots&\tilde{u}_{N,N}\end{array}\right)$ (11)
$\xrightarrow[]{}\left(\begin{array}[]{cccc}\tilde{u}_{1,1}&\tilde{u}_{1,2}&\cdots&\tilde{u}_{1,N}\\ 0&\tilde{\tilde{u}}_{2,2}&\cdots&\tilde{\tilde{u}}_{2,N}\\ \vdots&\vdots&\ddots&\vdots\\ 0&\tilde{\tilde{u}}_{N-1,2}&\cdots&\tilde{\tilde{u}}_{N-1,N}\\ 0&\tilde{u}_{N,2}&\cdots&\tilde{u}_{N,N}\end{array}\right)\xrightarrow[]{}\left(\begin{array}[]{cccc}\tilde{u}_{1,1}&\tilde{u}_{1,2}&\cdots&\tilde{u}_{1,N}\\ 0&\tilde{\tilde{\tilde{u}}}_{2,2}&\cdots&\tilde{\tilde{\tilde{u}}}_{2,N}\\ 0&0&\ddots&\vdots\\ 0&0&\cdots&\tilde{\tilde{\tilde{u}}}_{N-1,N}\\ 0&0&\cdots&\tilde{\tilde{\tilde{u}}}_{N,N}\end{array}\right)$ (22)
At the end all subdiagonal elements are zeroed out. Since the product of unitary matrices is still unitary, and the GR are unitary, the final matrix in eq. 11 is unitary and thus we deduce that coefficients above the diagonal are also null, leaving only diagonal elements of norm one. Taking $N_{GR}$ operations to obtain this result by applying succesive GR numbered by $i$, we end up in this situation.
$\left(\prod_{i=0}^{N_{GR}}GR_{i}\right)U$ $=\begin{bmatrix}e^{-i\xi_{1}}&0&\cdots&0\\ 0&\ddots&\ddots&\vdots\\ \vdots&\ddots&\ddots&0\\ 0&\cdots&0&e^{-i\xi_{N}}\end{bmatrix}$ (23)
We can make the z rotations needed to remove the $e^{-i\xi_{i}}$ using some combinations of x and y rotations, to end up with the identity matrix at the right of eq. 23, which means the applied operations would have constructed $U^{\dagger}$ (up to a global phase). Thus by inputing the conjugate transpose of a target unitary matrix $U$ in the GRD, we can construct $U$. An important point of eq. 5 is that, using the Hamiltonian in eq. 4 to make a GR, the amplitude $\Omega$ of a pulse $e^{-iHt}$ and its duration $t$ are equivalent and represented by the angle $\phi$. Thus a longer pulse of a smaller amplitude is equivalent to a shorter pulse of a larger amplitude as long as their product are equal. This would not be the case if we were to consider a drift term. In practice the power output of the experimental devices must not exceed a certain threshold, otherwise the pulses could damage the molecule or induce undesired transitions between energy levels. [10]. Moreover, mathematically, a higher amplitude pulse would induce a non-negligible Bloch-Siegert effect [5] due to the fast-oscillating terms neglected in the RWA, which would make eq. 4 not valid anymore. Thus it is not possible to make arbitrarily faster quantum gates just by considering pulses of larger amplitudes.
The GRD algorithm relies only on linear algebra, is interpretable and terminates in a finite number of steps. However these steps must be applied sequentially. On the other hand, the techniques relying on optimization generally do not terminate in a finite number of steps and could be attracted to local extrema, but they are able to handle multi-chromatic pulses and thereby achieve significantly faster implementations of gates.
II.1.2 An optimization algorithm: the GRadient Ascent Pulse Engineering (GRAPE) method
The GRAPE method [24] decomposes a quantum operation into a sequence of multi-chromatic pulses and follows a gradient descent described in [24, Section 2]. For numerical results in section III.2 we will use the GRAPE method from the QuTiP library since this package is widely used [21]. Assume one can model a physical system by an Hamiltonian composed of available control terms $H_{k}$ with $k\leq N_{\text{ctrls}}\in\mathbb{N}$ controlled by some continuous functions $u_{k}(t)$. We want to reach an objective gate $O$ by constructing a gate $U$ as close as possible and defined by (considering $\hbar=1$):
$H(t)$ $=\sum_{k=1}^{N_{\text{ctrls}}}u_{k}(t)H_{k}.$ (24)
$U$ $=e^{-i\int_{t=0}^{T}H(t)tdt}$
Since determining optimal controls in closed-form is usually out of reach, we choose a discretization scheme and numerically solve an non-linear program (direct approach). In this case, control functions are replaced by piecewise constant functions: for each piece, the duration is fixed to $\Delta t$ but the amplitudes are optimization variables. $O$ has to be reached with a finite sequence of $N$ gates $U_{j}$ with $j\leq N\in\mathbb{N}$ that write as:
$U_{j}$ $=e^{-i\Delta t(H_{0}+\sum_{k=1}^{N_{\text{ctrls}}}u_{j,k}H_{k})},$
$U$ $=\prod_{j=1}^{N}U_{j}.$
In a Hilbert space of dimension $d$, the cost function to be minimized reflects the error between the U and the target gate, up to a global phase. With $\mathbf{u}=\{u_{j,k}\}_{j=1,\dots,N}^{k=1,\dots,N_{\text{ctrls}}}$. It is given by
$\text{Cost}(\mathbf{u})=1-\frac{1}{d}\left|\operatorname{Tr}(O^{\dagger}U)\right|.$ (25)
and referred to as infidelity, following [21]. To update the control terms $u_{j,k}$ different optimization algorithms could be used. QuTiP [21] uses the L-BFGS-B [9] method of SciPy [35] for instance.
II.1.3 Comparing execution times between methods
To ensure a fair comparison of the execution times obtained with the methods considered in this work, the sum of the amplitude of each monochromatic pulses is bounded by a fixed amplitude $\Omega$. This is inspired from the analytical GRD approach, where the control consists of monochromatic pulses with fixed amplitude $\Omega$, chosen to be the maximum value allowed by the experimental setup, and variable duration. More precisely, once $\theta$ and $\phi$ are determined from eq. 5, the pulse duration $t$ is chosen such that $\phi=\Omega t$. By contrast, optimization methods such as the GRAPE method employ discretized multichromatic pulses defined over fixed time intervals, with independently varying amplitudes that may exceed the maximum value achievable by the arbitrary waveform generator. Therefore, in order to compare the optimal-control methods with the analytical GRD protocol on equal footing, it is necessary to recast the Hamiltonian applied during each discretized time step $\Delta t$ into an equivalent form satisfying the same amplitude bound. More specifically, for a frequency component indexed by $i$, with control operators $\sigma_{x,i}$ and $\sigma_{y,i}$, and with corresponding real-valued amplitudes $\alpha_{i}$ and $\beta_{i}$, the Hamiltonian during a time step $\Delta t$ is transformed as follows:
$H$ $=\sum_{i=1}^{N_{\text{ctrls}}}\left[\alpha_{i}\sigma_{x,i}+\beta_{i}\sigma_{y,i}\right]$
$=\sum_{i=1}^{N_{\text{ctrls}}}\left[\alpha_{i}\sigma_{x,i}+\beta_{i}\sigma_{y,i}\right]=\Delta t\sum_{i=1}^{N_{\text{ctrls}}}\Omega_{i}\left[\cos(\theta_{i})\sigma_{x,i}+\sin(\theta_{i})\sigma_{y,i}\right].$ (26)
With
$\Delta t$ $=\sum_{j=1}^{N_{\text{ctrls}}}\Omega_{j},$ (27)
$\Omega_{i}$ $=\frac{\sqrt{\alpha_{i}^{2}+\beta_{i}^{2}}}{\Delta t},$ (28)
$\theta_{i}$ $=\text{atan2}\left(\frac{\beta_{i}}{\alpha_{i}}\right).$ (29)
We have to normalize the $\Omega_{i}$ for the sum of the $\Omega_{i}$, representing the energy of the pulses, to not go higher than $\Omega$ for experimental reasons as explained at the end of section II.1.1. We take 1 to be the $\Omega$ maximum value to make the link with a similar scaling of $\phi$ in eq. 5. This framework means that it takes $\Omega^{-1}\times\pi/2$ units of time to inverse two populations, with this $\Omega$ being the maximum available amplitude sustainable by the system. The comparison we propose is based on the equivalence between shorter durations and higher amplitudes, or vice-versa, and as such is relevant only when the drift Hamiltonian can be neglected.
II.2 Presenting a shooting technique: the Method for Adjoint and Gradient-based self-Iterative Construction And Refinement of Pulses (MAGICARP)
hThe Method for Adjoint and Gradient-based self-Iterative Construction And Refinement of Pulses (MAGICARP) is a method [18] and named as such to distinguish it from conventional gradient techniques. The core of this article lies in its numerical implementation for a system inspired from the physics of SMMs, especially for the triple and double deckers systems detailed in section III.1.
II.2.1 Pontryagin’s Maximum Principle (PMP)
The MAGICARP consists in a finite-dimensional parametrization of controls in feedback form which is derived from the Pontryagin maximum principle (PMP), a mathematical tool to reframe an optimisation problem on “trajectories” (living in an infinite dimensional space) into an Hamiltonian problem where only a finite amount of coefficients, encoding the trajectories, are to be optimized. This change of point of view is analogous to using a principle of least action in classical physics. The problem of optimisation of the control fields can be rewritten as a problem where
$\forall t,\,\dot{U}(t)$ $=F(U(t),u(t),t),$
$U(0)$ $=U_{0},$
$U(t_{f})$ $=U_{f},$
$\forall t,\,u(t)$ $\in\mathbb{R}^{N_{\text{ctrls}}}$
with $u(t)$ a vector of all the $u_{k}(t)$ for $k\leq N_{\text{ctrls}}\in\mathbb{N}$, $U(t)$ the gate obtained at time $t$ via the optimisation protocol, $U_{0}$ that is usually the identity and $U_{f}$ the target gate. To select a control among all those steering the system between the desired states, we seek to minimize a cost function $\mathcal{C}$ including a terminal cost $G$ and a running cost $F_{0}$ reflecting what a desirable control is:
$\mathcal{C}=G(U(t_{f}),t_{f})+\int_{0}^{t_{f}}F_{0}(U(t),u(t),t)\,dt.$ (30)
To achieve the transfer as quickly as possible, $G$ could either be the duration $t_{f}$ or, if we drop the hard constraint $U(t_{f})=U_{f}$, the infidelity between $U(t_{f})$ and $U_{f}$. In $F_{0}$ we could include the fact that the pulse amplitude has to be maximized with $F_{0}=-\sqrt{\sum_{k}{u_{k}(t)}^{2}}$ for instance. If the pulse amplitude ends up being higher than $\Omega$ the maximum amplitude of the pulses, we could always renormalize the controls by applying them for a longer time like we did for GRAPE thanks to the equivalence between amplitude and time explained in section II.1.1 and detailed in section II.1.3.
The PMP [32] is an important tool of optimal control theory. To strengthen the physical intuition behind it, we provide a heuristic derivation inspired from [3] of the corresponding necessary conditions of optimality by connecting it to the Euler-Lagrange equations from the calculus of variations. Starting from static initial and final target quantum gates (That is, static boundary conditions), it turns the problem of optimizing a cost function into the problem of minimizing an action
Definition II.1.
Let $U(t)\in\mathbb{R}^{n\times n}$ be the quantum gate we are constructing, at time $t$, $\Lambda(t)\in\mathbb{R}^{n\times n}$ an adjoint state, defining the optimisation we choose for the problem., and $u(t)\subset\mathbb{R}^{N_{\rm ctrls}}$ the controls $u_{k}(t)$ with $k\leq N_{\rm ctrls}\in\mathbb{N}$ in eq. 24. Let $G$ be a terminal cost function, for instance the fidelity, $F_{0}$ a running cost and $F$ a vector function describing the system dynamics. The optimal control is defined as the one minimizing the action $S$:
$S$ $=\int_{0}^{t_{f}}dt\mathcal{L}+G(U(t_{f}),t_{f}),$
$\mathcal{L}$ $=\left[F_{0}(U(t),u(t),t)+\Lambda(t)\left(\dot{U}(t)-F(U(t),u(t),t)\right)\right]$
$\Lambda$ $=\frac{\partial\mathcal{L}}{\partial\dot{U}}$
The function under the integral is the Lagrangian $\mathcal{L}$ of the optimal control problem and $\Lambda(t)$ is an adjoint state. The optimal control is then defined as a minimum of the action $S$. The first variation $\delta S$ of the action in response to a perturbation $\delta u$ of the control is formally given in terms of the state variation $\delta U$ by
$\delta S$ $=\int_{0}^{t_{f}}\left[\frac{\partial F_{0}}{\partial U}\delta U+\frac{\partial F_{0}}{\partial u}\delta u+\delta\Lambda(\dot{U}-F)+\Lambda\left(\delta\dot{U}-\frac{\partial F}{\partial U}\delta U-\frac{\partial F}{\partial u}\delta u\right)\right]+\frac{\partial G}{\partial U(t_{f})}\delta U(t_{f}),$
$\delta S$ $=\int_{0}^{t_{f}}dt\Bigg(\left[\frac{\partial F_{0}}{\partial U}-\dot{\Lambda}-\Lambda\frac{\partial F}{\partial U}\right]\delta U+\left[\dot{U}-F\right]\delta\Lambda+\left[\frac{\partial F_{0}}{\partial u}-\Lambda\frac{\partial F}{\partial u}\right]\delta u\Bigg)$
$+\left(\frac{\partial G}{\partial U(t_{f})}+\Lambda(t_{f})\right)\delta U(t_{f})+\Lambda(0)\delta U(0),$
by integration by parts of the term $\Lambda\delta\dot{U}$. Thus the minimums of $S$ fulfill the following conditions
Proposition II.2.
If a control $u$ associated to the state $U$ is optimal, then there exists an adjoint state $\Lambda$ such that the following conditions hold:
$\dot{\Lambda}$ $=\frac{\partial F_{0}}{\partial U}-\Lambda\frac{\partial F}{\partial U},$ (31)
$\dot{U}$ $=F,$ (32)
$\frac{\partial F_{0}}{\partial u}$ $=\Lambda\frac{\partial F}{\partial u},$ (33)
$U(0)=U_{0}$ $,\quad U(t_{f})=U_{f}.$
From a physical point of view, eq. 31 describes an Euler-Lagrange equation for $U$, eq. 32 describes the dynamics of the system (an Euler-Lagrange equation for $\Lambda$) and eq. 33 imposes a condition on the controls (a Euler-Lagrange equation for $u$). It is then possible to translate this formulation into an Hamiltonian problem, with the Pontryagin Hamiltonian defined as:
$H_{P}$ $=\Lambda\cdot\dot{U}-\mathcal{L}=\Lambda\cdot F-F_{0}.$
With $\cdot$ denoting the Hilbert–Schmidt inner product. Given that we will follow a physcial trajectory of the system where eq. 32 is satisfied, the solution to the optimization problem is then written as
$H_{P}$ $=\Lambda\cdot F-F_{0},$ (34)
$\dot{\Lambda}$ $=-\frac{\partial H_{P}}{\partial U},$ (35)
$\dot{U}$ $=\frac{\partial H_{P}}{\partial\Lambda},$ (36)
$\vec{0}$ $=\frac{\partial H_{P}}{\partial u},$ (37)
$U(0)=U_{0}$ $,\quad U(t_{f})=U_{f}.$ (38)
II.2.2 The MAGICARP
For a quantum evolution the eq. 32 corresponds to the Schrödinger equation and $U_{0}$ in eq. 38 is the identity. Inputing the running cost $F_{0}=-\sqrt{\sum_{k}{u_{k}^{2}(t)}}$ in the Hamiltonian optimization problem of eq. 34 Jankovic derived in [18]
Proposition II.3.
Pulses $u_{k}(t)$ of eq. 24 with $k\leq N_{\rm ctrls}\in\mathbb{N}$ that minimize the running cost $F_{0}$ are of the form
$u_{k}(t)=\operatorname{Re}\operatorname{Tr}\left(U(t)MU^{\dagger}(t)H_{k}\right),$ (39)
$U(t)$ is the time-evolution operator of the controlled system at time $t$ and $M$ is a constant, traceless and Hermitian matrix.
Proof.
Inputing $F_{0}=-\sqrt{\sum_{k}{u_{k}(t)}^{2}}$ in eq. 34, and using eq. 36 with the usual expression of $\dot{U}$ as iHU-iHU with $H$ the Hamiltonian of the system (in eq. 24) gives
$H_{P}$ $=\operatorname{Re}\operatorname{Tr}\left(-i\Lambda^{\dagger}HU\right)+\sqrt{\sum_{k}{u_{k}(t)}^{2}}.$ (40)
So we can replace in eq. 35 and eq. 37 to obtain the following equations:
$H_{P}$ $=\operatorname{Re}\operatorname{Tr}\left(-i\Lambda^{\dagger}HU\right)+\sqrt{\sum_{k}{u_{k}(t)}^{2}},$ (41)
$\dot{\Lambda}$ $=-iH\Lambda,$
$\dot{U}$ $=-iHU,$ (42)
$\vec{0}$ $=\operatorname{Re}\operatorname{Tr}\left(-i\Lambda^{\dagger}\frac{\partial H}{\partial\vec{u}}U\right)+\frac{\vec{u}}{\sqrt{\sum_{k}u_{k}^{2}}},$ (43)
$U(0)=U_{0}$ $,\quad U(t_{f})=U_{f}$
Then a very useful proposition to notice is that eq. 42 is easy to solve. By differentiating the function $\Lambda(t)^{\dagger}U(t)$, we find that its derivative is everywhere zero. Hence $\Lambda(t)=U(t)\Lambda(0)$. Since $\Lambda(0)$ is in the Lie algebra of the special unitary group, $\Lambda_{0}=iM$ for some traceless Hermitian matrix $M$. Then we can rewrite $i\Lambda^{\dagger}$ in eq. 43 as $MU^{\dagger}$ with $M$ a constant, traceless, Hermitian matrix. Finally by using the cyclic proposition of the trace and $\frac{\partial H}{\partial u_{k}}=H_{k}$ for all $k$ we obtain
$\frac{u_{k}(t)}{c(t)}=\operatorname{Re}\operatorname{Tr}\left(U(t)MU^{\dagger}(t)H_{k}\right),$
where
$c(t)=\sqrt{\sum_{k}u_{k}^{2}(t)},$
We could see the c(t) as constant in time if we maximize the amplitude of the controls by this quantity when there is no drift, so we end up with the expression of the optimal controls given in eq. 39. ∎
MAGICARP is a parametrization of controls by means of the initial matrix $M$. By means of a suitable choice of optimization routine, starting from an initial guess, this parameter is iteratively updated until the target gate is attained up to some user-defined tolerance. As such, it belongs to the class of shooting methods in which, roughly, one aims at finding the appropriate initial covector in such a way that the optimal set of equations derived from the PMP are satisfied while the desired target state is attained with sufficient precision. In particular MAGICARP suffers from the necessity of finding relevant initial guesses for the parameter $M$. However what is worth noting here is that, by working with the propagator or flow of the equation of motion and exploiting the Lie group structure, the maximization condition of the PMP leads to a control in closed-loop form. In other words, the control amplitudes are determined only by the instantaneous propagator $U(t)$. One notable advantage is that the parameter $M$ is not tied to any choice of discretization scheme and fully determines the associated control given the control Hamiltonians.
Remark 1.
Proposition II.3 is only concerned with the so-called normal extremals of the Pontryagin maximum principle. Put differently, for some choices of control Hamiltonians and some target gates, the optimal controls might not be given by Equation (39) for some traceless Hermitian matrix $M$. Nevertheless it appears natural to exploit numerically this simple parametrization in an attempt to obtain more time-efficient controls pulses. See Remark 2 in Appendix C for a discussion.
III Numerical implementation of the MAGICARP on Single Molecule Magnets (SMM)
III.1 The physics of SMM
The concept of Single-Molecule Magnets used as quantum devices started in the 90s [36, 23]. Their interest lies in their very small size and the fact that they show quantum effects. We will compare numerical results obtained via the MAGICARP on these two molecules.
III.1.1 SMM
The first molecule of interest is the bisphthalocyaninato terbium(III) ($\text{Tb}\text{Pc}_{2}$) molecule, a SMM synthesised in 2003 [17] and shown in the fig. 1a which is also known as the double decker. The $\text{Tb}^{3+}$ ion has a nuclear spin of $I=3/2$ [14, 19], which gives rise to $2I+1=4$ energy levels. This makes the $\text{Tb}\text{Pc}_{2}$ an effective qu-4-it, a qudit [37] with $d=4$. It is also possible to superpose two double deckers into a triple decker $\text{Tb}_{2}\text{Pc}_{3}$ molecule, which then acts as an effective qu-16-it [4] thanks to a coupling between the two $\text{Tb}^{3+}$ ions [22].
Figure 1: (a) A $\text{Tb}\text{Pc}_{2}$ double decker, composed of the central $\text{Tb}_{3+}$ ion (red) surrounded by nitrogen (blue) and carbon (grey) atoms of the phthalocynanine ligands. Image from [16]. (b) A triple decker, $\text{Tb}_{2}\text{Pc}_{3}$. The two dark blue atoms are also $\text{Tb}^{3+}$ ions. Image from [16].
To construct quantum gates we need to induce transitions between the system’s energy levels by applying an electromagnetic pulse with an appropriate frequency. Not all transitions are allowed. In the example of the double decker, selection rules only allows for addressing transitions between levels separated by a $\Delta I=\pm 1$ [19].
Figure 2: (a) The linear graph connecting the four energy levels of a double decker $\text{Tb}\text{Pc}_{2}$. The spacing is unequal thanks to a non linear term in the Hamiltonian. Otherwise energy levels would be positioned on the dashed lines. Values from [14]. (b) The graph connecting the sixteen energy levels of a triple decker $\text{Tb}_{2}\text{Pc}_{3}$. Each node represent an energy level, ranked by their energy. The edges represent addressable Rabi transitions between nodes, for a certain energy difference in Kelvin. The energy levels are denoted by “lvl” with the first digit denoting the nuclear energy level for the first molecular magnet and the second digit the other nuclear energy level. This figure is done using parameters that differ slightly from [4], and at a large magnetic field to suppress level mixing.
This means that only transitions shown in fig. 2a have an effect on the system. For instance, an energy level corresponding to I=1/2I=-1/2 can communicate when excited at the right frequency with the energy levels corresponding to I=3/2I=-3/2 or $I=1/2$ but not with the level corresponding to $I=3/2$. Thus only neighbouring interactions are allowed, resulting in what we refer to as a linear graph. The resonances can be individually adressed thanks to a non-linear term that results from an electric quadrupolar interaction. For the case of the triple decker, only coupling between adjacent levels of nuclear spin $I_{1}$ and $I_{2}$ ($\Delta I=1$) are allowed, the 1 or 2 indices denoting the nuclear energy levels for each of the $\text{Tb}^{3+}$ ions. The non trivial resulting graph is shown in fig. 2b. Usually the allowed transitions are determined experimentally and as long as all energy levels are connected, it is possible to create any unitary gate on the Hilbert space associated to these SMM, for instance by using the GRD.
III.1.2 Running a Quantum Fourier Transform (QFT) on a SMM
[…]
Figure 3: Control amplitudes (left panels) and population dynamics (right panels) for the implementation of a QFT gate on the double decker graph. Panels (a)–(b) correspond to GRD, (c)–(d) to GRAPE, and (e)–(f) to MAGICARP. In (a), each blue activation, delimited by black lines, correspond to a different $(\theta,\phi)$ for a GR in eq. 5. The length of each blue activation is $\phi$. For GRAPE and MAGICARP, each color denotes a distinct control amplitude. The target fidelity is $10^{-4}$. GRAPE controls are discretized in 50 time steps. All times are given in units of $\Omega^{-1}$ (see section II.1.3).
[…]
III.2 Numerical results
For both the GRAPE and MAGICARP methods, we set the target infidelity to $10^{-4}$ in this work. We explain in section A.2 why this is a convenient choice. The MAGICARP allows optimization of only a finite number of parameters (the constant coefficients of the matrix $M$ in eq. 39) rather than the infinite dimensional space of the control $u_{k}(t)$, that are smooth functions of time. An interesting proposition of this method is that the controls obtained fullfill a necessary (albeit not sufficient) condition for a time-optimal control.
Figure 4: Comparison of GRAPE and MAGICARP execution times to implement 1000 random Haar unitaries. (a) For a $\text{Tb}\text{Pc}_{2}$ molecule. (b) For a $\text{Tb}_{2}\text{Pc}_{3}$ molecule.
It appears numerically that this parametrization yields shorter gate times, all the more as the dimension of the Hilbert space increases. For the triple decker in fig. 4b the smallest execution time found by the MAGICARP is 37 % faster than for the GRAPE method for instance. An important point is that it is increasingly computationally difficult to optimize the $M$ matrix of eq. 39 with the growing size of the Hilbert space. In appendix C, we discuss an optimizer based on the natural gradient which provides solutions below the infidelity threshold for qudits up to $d=16$ dimensions, e.g. the triple decker discussed in Section Y. We used it to obtain the results of fig. 4b. A Julia and a Python code are available on the Github repositories Magicarp and pyMagicarp. Overall, the higher the dimension of the system is, the better the results of the MAGICARP in comparison to GRAPE when comparing median execution times for random gates or the minimum execution time for a specific gate (the QFT) as shown in fig. 5.
Figure 5: Comparison of GRAPE and MAGICARP execution times on linear graphs with 3 to 9 nodes, as a function of the Hilbert space dimension. There is at least 100 initializations for each dimension. We used the natural gradient descent described in appendix C for the green curves, and for systems of dimension below 6, we also used the more computationally expensive L-BFGS-B optimizer from Scipy [35] described by the yellow curve. The ratios indicate how much faster MAGICARP is compared to GRAPE for a given dimension. (a) Ratio of the median execution times for random Haar gates using GRAPE with random initializations over the median time using MAGICARP. (b) Ratio of the minimum execution times for a QFT gate using GRAPE with random initializations over the minimum time using MAGICARP.
We choose to use a QFT gate to compare the minimum execution time for the MAGICARP and the GRAPE method since this gate is not trivial, is key to many important quantum algorithms as explained in section III.1.2, and has execution times similar to a random gate (gaussian-like) as one can see when comparing fig. 4b and fig. 6a. Note that for small dimensions (3 and 4), their is no improvement of the execution times with the natural gradient optimizer used here since the ratios are below 1 as shown in fig. 5 for the corresponding curve. Though if we compare the MAGICARP with the same gradient method as GRAPE (A L-BFGS-B gradient from SciPy [9, 35].), with the same parameters, and using automatic differentiation, we can compute the results at low dimensions (up to 5) and do witness that the MAGICARP is at least faster than GRAPE both in median time for random gates and in minimum time for the QFT gate for this naive implementation. This observation illustrates that the advantage of MAGICARP over GRAPE discussed in this paper leaves room for further improvements of the execution time by carefully designing the algorithm used to optimize the parameter of MAGICARP.
Comparing these methods only by implementing target gates chosen uniformly at random in the unitary group provides insights into some kind of improvement of the execution time “on average”. However, in practice, not all unitary matrices are of interest for the existing quantum algorithms. We therefore compare the GRAPE and MAGICARP methods on the triple decker ($d=16$) by implementing four particular commonly used quantum gates: QFT, T, X and SUMX [37]. The latter generalizes the CNOT gate to qudits. We obtained significantly better results in fig. 6 for the MAGICARP, regarding both median and minimum execution times.
Figure 6: Comparison of GRAPE and MAGICARP execution times for 1000 random initializations on a $\text{Tb}_{2}\text{Pc}_{3}$ molecule. (a) QFT gate. (b) T gate. (c) X gate. (d) SUMX gate.
IV Conclusion
This work deals with the numerical implementation of a proof of concept to design smooth control pulses using a parametrization derived from the Pontryagin maximum principle. The corresponding method, coined MAGICARP, is applied to the implementation of quantum gates on spin qudit systems while negelecting the drift term. We demonstrated that, relative to the GRAPE method, the MAGICARP can yield significantly better execution times for spin systems up to dimension sixteen, and all the more as the dimension of the system is large. The method was showcased on two single molecule magnets, the $\text{TbPc}_{2}$ and $\text{Tb}_{2}\text{Pc}_{3}$ molecules, taking into account some of their specific constraints.
Several promising avenues for future research are worth pursuing:
  1. 1.
    Experimentally test the effectiveness of the pulses derived from the MAGICARP.
  2. 2.
    Extend the MAGICARP parametrization to closed system involving a drift Hamiltonian, and open systems.
  3. 3.
    Understand which optimizations and parametrizations could power up the MAGICARP, for instance by finding a clever gradient descents algorithm.
  4. 4.
    Relate the connectivity of the graph between the system’s energy levels to the efficiency of the method in terms of execution times.
Acknowledgements
The authors would like to express their gratitude to Benjamin Bakri for their fruitful discussions. The authors would like to acknowledge the High Performance Computing Center of the University of Strasbourg for supporting this work by providing scientific support and access to computing resources. Part of the computing resources were funded by the Equipex Equip@Meso project (Programme Investissements d’Avenir) and the CPER Alsacalcul/Big Data. This work of the Interdisciplinary Thematic Institute QMat, as part of the ITI 2021-2028 program of the University of Strasbourg, CNRS and Inserm, was supported by IdEx Unistra (ANR-10-IDEX-0002), and by SFRI-STRAT’US project (ANR 20 SIFRI 0012) and EUR QMAT (QMAT ANR-17-EURE-0024) under the framework of the French Investments for the Future Program. This work was carried out as part of the Inria Exploratory Action MARCQ, supported by Groupe La Poste, a patron of the Inria Foundation.
Appendix A MAGICARP
A.1 Importance of the graph
The graph on which simulations are taken are ones like the double decker from fig. 2a, linearly coupled graphs, and the one of the triple decker from fig. 2b. If we assume different graphs, possible in other physical molecules and platforms, we could have different time implementations for the same target gates and Hilbert space dimension. A most simple configuration is the linear graph where each level is only coupled to its nearest neighbours. It’s the congifuration of the fig. 1a. Another extremal configuration is the complete graph where each level is coupled to all the others. We can test how the GRAPE method and the MAGICARP take advantage of the growing connectivity of the graph. We will plot how the MAGICARP and the GRAPE method compare to themselves when runned over a complete graph rather than on a linearly coupled one in fig. 7. Note how in both case of fig. 7b the smallest execution time found is increasingly improved by the fact of using a complete graph instead of a linear one (thus having more available controls) with the size of the Hilbert space. In the case of fig. 7a the MAGICARP is slightly helped by the use of a complete graph instead of a linear one (gates are in median around 20% faster in all dimension) but is not as much of an improvement as for GRAPE. Thus sparsely connected graph seem already convenient for the MAGICARP.
Figure 7: Figure comparing the median or the minimum time to make at least 100 gates with a precision of 1e-4 for the MAGICARP and the GRAPE method on complete graphs with 3 to 9 nodes with themselves on linear graphs. For each optimization method, the quantity shown is the ratio between the runtime on a linear graph and the runtime on a complete graph, effectively describing how each method is faster when being run on a complete graph instead of a linear one. (a) Ratio of the median execution times for 100 random Haar gates. (b) Ratio of the minimum execution times for a QFT gate.
A.2 Importance of a $\sigma_{z}$ control
In all previous simulations we were using a control on $\sigma_{x}$ and $\sigma_{y}$ for each edge between levels of the Hilbert space, like in eq. 4, but not on $\sigma_{z}$ since usual physical platforms (especially for SMMs) are not allowing for such a control for experimental reasons. If we lift this experimental constraint, we can test how the GRAPE method and the MAGICARP take advantage of having a control on $\sigma_{z}$ on different graphs. We will plot how the MAGICARP and the GRAPE method compare to themselves when runned while allowing to play on this $\sigma_{z}$ control rather than only $\sigma_{x}$ and $\sigma_{y}$ for each edge between levels of a linear and a coupled graph in fig. 8.
Figure 8: Figure comparing the median or the minimum time to make at least 100 gates with a precision of 1e-4 for the MAGICARP and the GRAPE method on complete and linear graphs with 3 to 9 nodes with themselves when allowing for a control on $\sigma_{z}$. For each optimization method, the quantity shown is the ratio between the runtime with only $\sigma_{x}$ and $\sigma_{y}$ controls and the runtime with all three controls on each edge of the graph the method has been run on. It effectively describes how each situation (method + graph it is run on) is faster when being run while allowing a control on $\sigma_{z}$ instead of just $\sigma_{x}$ and $\sigma_{y}$. The red curves correspond to the aforementioned ratio for the GRAPE method run on a linear graph, the orange ones for the GRAPE method run on a complete graph, the green ones for the MAGICARP run on a lienar graph, and the cyan ones for the MAGICARP run on a complete graph. The x-axis describes the Hilbert space dimension the optimization methods have been run on. (a) Ratio of the median execution times for 100 random Haar gates. (b) Ratio of the minimum execution times for a QFT gate.
One can note that adding a control on $\sigma_{z}$ is not improving much the median solutions found on random gates in fig. 8a or the minimum solutions found in fig. 8b for the specific QFT gate in all situations but in the case of the MAGICARP method run on a complete graph, the cyan curve of fig. 8a, especially when the dimension grows. It applies only for the median execution time since the same cyan curve in fig. 8a is not affected. It would be interesting to understand further why adding a $\sigma_{z}$ control gives a particular advantage to the MAGICARP on a strongly connected graph.
Appendix B Benchmark
B.1 Improvement of the GRAPE with the target fidelity
The GRAPE method gives better execution times according to the target fidelity it has too reach.
Figure 9: Figure comparing the median and minimum time found for making the same 1000 random Haar gates for the GRAPE with varying target fidelities. (a) Figure for a Hilbert space dimension of 4 on the double decker graph. (b) Figure for a Hilbert space dimension of 16 on the triple decker graph.
It is understandable since being able to be further away from the target obviously relieves a constraint on the pulses. We chose to compare execution times at 1e-4 of target fidelity for the GRAPE when comparing to other algorithms for multiple reasons. First, this is already a fidelity allowing for making a short algorithm without being bothered by this lack of precision of the gates while in many platforms noise is already inducing an error of this order [28]. Second, Maintaining the fidelities of pulses found by the MAGICARP to be strictly lower than 1e-4 proved difficult with the gradient we used. Third, at lower fidelities than 1e-4 GRAPE is not giving significantly slower pulses than for this value of the target fidelity as shown infig. 9.
B.2 Improvement of the GRAPE with the number of controls
The GRAPE method gives better execution times according to the number of piece-wise amplitude steps it can use, though up to a limit. In all this article, we chose to compare execution times with $10(\text{dim}+1)$ piece-wise amplitude steps for each controls, with dim being the dimension of the Hilbert space at hand. It is a large enough number of amplitude to not see improvements by taking a larger amount of piece-wise amplitude steps for each controls as shown infig. 10.
Figure 10: Figure comparing the median and minimum time found for making the same 1000 random Haar gates for the GRAPE with varying number of piece-wise amplitude steps for each controls. The x-axis describe a multiplicator $x$ to the number of controls used, this number being $x(\text{dim}+1)$ with dim the dimension of the Hilbert space. The target fidelity is 1e-4. (a) Figure for a Hilbert space dimension of 4 on the double decker graph. (b) Figure for a Hilbert space dimension of 16 on the triple decker graph.
Appendix C Numerical framework
Objectives.
Our objectives in this section is twofold. We provide closed-form expressions for the derivative of the map that associates to every control parameter $g$ the final propagator of Schrödinger’s equation. We outline its derivation which relies on the so-called adjoint state method. This provide the tools needed to implement any first-order optimization routine and possibly to adapt the $M$ parametrization of controls to another problem. On the other hand, we propose our implementation of the natural gradient descent, an optimization routine used to produce some of the figures in this paper. Julia and Python codes are available on the Github repositories Magicarp and pyMagicarp.
C.1 Two equivalent optimal control problems in the driftless case
Let $\operatorname{i}\mathfrak{su}(d)$ denote the set of traceless Hermitian matrices of order $d$.
Assumption C.1 (Target).
$U_{f}\in SU(d),\quad U_{f}\neq I$
Assumption C.2 (Orthonormal control Hamiltonians).
$H_{j}\in\operatorname{i}\mathfrak{su}(d),\quad\operatorname{Re}\operatorname{Tr}(H_{i}H_{j})=\delta_{i,j},\quad 1\leq i,j\leq m$
Assumption C.3 (Lie algebra rank condition).
$\operatorname{Lie}(\operatorname{i}H_{1},\ldots,\operatorname{i}H_{m})=\mathfrak{su}(d).$
Let us comment on the second and third assumptions. In some applications the control Hamiltonians are related to the Pauli matrices which are orthogonal as in Assumption C.2. As such this assumption is not too restrictive. Together with the normalization condition, it is convenient but by no means necessary.
The third Assumption C.3 is common in the practical coherent control of quantum systems because the control Hamiltonians often do not linearly span the set of tangent vectors to the special unitary group at the identity matrix. The Lie algebra rank condition is about iterated commutators of the control Hamiltonians. When it holds, combining the available control directions provides a family of tangent vectors that is sufficiently rich to move in any admissible direction. Intuitively this is because to second order in $t$ close to zero, $e^{-tA}e^{-tB}e^{tA}e^{tB}=I+t^{2}[A,B]+o(t^{2})$. All in all Assumption C.3 is meant to guarantee that any desired special unitary gate is reachable by means of an appropriate sequence of constant pulses.
Two optimal control problems.
Let $P$ stand for the orthogonal projector onto the real linear span of $H_{1},\ldots,H_{m}$ for the Frobenius (Hilbert-Schmidt) norm $M\mapsto\operatorname{Tr}(MM^{\dagger})^{1/2}$. Consider the following two optimal control problems. On the one hand
$\inf_{u\in L^{\infty}(0,T;\mathbb{R}^{m})}T$ ($\mathscr{P}_{u}$)
subject to the pointwise constraint $\sum_{j=1}^{m}u_{j}(t)^{2}\leq 1$ and to the terminal constraint $U(T)=U_{f}$, where the propagator $U$ solves the Cauchy problem $U(0)=I$ and $\dot{U}(t)=-\operatorname{i}\sum_{j=1}^{m}u_{j}(t)H_{j}U(t)$ for $0<t<T$.
On the other hand
$\inf_{M\in\operatorname{i}\mathfrak{su}(d)}|P(M)|$ ($\mathscr{P}_{M}$)
to the terminal constraint $U(1)=U_{f}$, where the propagator $U$ solves the Cauchy problem $U(0)=I$ and
$\dot{U}(s)=-\operatorname{i}\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}\left(U(s)^{\dagger}H_{j}U(s)M\right)H_{j}U(s),\quad 0<s<1.$ (44)
Using Assumption C.3 and rescaling, we can build an admissible control for Problem ($\mathscr{P}_{u}$) which steers the identity to $U_{f}$ while satisfying the experimental constraint $\sum_{j=1}^{m}u_{j}(t)^{2}\leq 1$. This relies on the Chow-Rashevskii theorem or more specifically [2, Corollary 3.1]. Moreover it can be shown that an optimal solution exists and that, using the maximization condition from Pontryagin’s maximum principle (see e.g. [8]), any normal extremal is admissible for Problem ($\mathscr{P}_{M}$) with
$T=|P(M)|,\quad\text{ and }\quad u_{j}(t)=\operatorname{Re}\operatorname{Tr}\left(U(s)^{\dagger}H_{j}U(s)\frac{M}{|P(M)|}\right),\quad\text{where}\quad\quad Ts=t.$ (45)
The time $T$ in the first problem corresponds in the second problem to the norm of the projection of $M$ onto the space of controllable directions, that is $|P(M)|$.
Remark 2 (Restriction to normal extremals).
Not all Pontryagin extremal are normal and hence, it is not true in general that every optimal control $u$ of Problem ($\mathscr{P}_{u}$) is of the form given by Equation (45). Depending on the choice of control Hamiltonians, one can prove that any abnormal extremal, which is not normal at the same time, cannot be optimal and hence can be disregarded. This occurs for instance in the $K+P$ setting. The proof given in [7, Appendix C] relies on the Cartan decomposition of Lie groups and the Goh necessary conditions derived in [1].
In Problem ($\mathscr{P}_{u}$) the constraint $|u(t)|\leq 1$ physically reflects an upper bound on the power transmitted to the qudit and, mathematically, such a constraint ensures the existence of an optimal control. Indeed without this constraint and if $U_{f}$ is non-trivial, given a control $u$ driven the system $U$ from $U(0)=I$ to $U(T)=U_{f}$ in time $T$, the more powerful control $u_{n}(t)=nu(t)$ performs the transfer in the shorter time $T_{n}=T/n$. Since $n$ can be arbitrarily large, the transfer can be performed in arbitrarily short time. However there is no control performing the transfer instantaneously with $T=0$ because this would imply $U(T)=U(0)$ but $U_{f}\neq I$ by assumption. In other words, the infimum is zero but is not attained and therefore not a minimum.
C.2 Differential of the end-point mapping and its adjoint operator
In the framework of Problem ($\mathscr{P}_{M}$), the state $U:[0,1]\to SU(d)$ solves a non-linear equation with feedback control determined by $M\in\operatorname{i}\mathfrak{su}(d)$. More precisely we are interested in the Cauchy problem
$U(0)=I,\quad\dot{U}$ $=-\operatorname{i}\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}(U^{\dagger}H_{j}UM)H_{j}U,\quad t\in(0,1).$ (46)
Let $E:\operatorname{i}\mathfrak{su}(d)\to SU(d)$ denote the end-point mapping, defined by $E(M)=U(1)$ where $U$ is the solution of (46) with parameter $M$. The derivative $E^{\prime}$ of this mapping is needed to implement a first order optimization routine in the so-called optimize-then-discretize framework. We now focus on computing this derivative.
Equip the space of traceless Hermitian matrices $\operatorname{i}\mathfrak{su}(d)$ and $\mathbb{C}^{d\times d}$ with the inner-product $\langle a,b\rangle=\operatorname{Re}\operatorname{Tr}(ab^{\dagger})$. Then at any point $M$, the derivative $E^{\prime}(M):\operatorname{i}\mathfrak{su}(d)\to\mathbb{C}^{d\times d}$ and its adjoint $E^{\prime}(M)^{\dagger}:\mathbb{C}^{d\times d}\to\operatorname{i}\mathfrak{su}(d)$ are obtained as follows. As a shorthand notation, we write $h_{j}(t)=U(t)^{\dagger}H_{j}U(t)$.
Lemma C.4 (Jacobian-vector product).
Given a perturbation $\delta M\in\operatorname{i}\mathfrak{su}(d)$, the resulting perturbation $\delta U(1)$ of the end-point $U(1)$ of the curve $U(t)$ is given by $E^{\prime}(M)\delta M=-\operatorname{i}U(1)w(1)$ where $w(t)\in\operatorname{i}\mathfrak{su}(d)$ satisfies
$w(0)=0,\quad\dot{w}=\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}\left(h_{j}\delta M-2\operatorname{i}Mh_{j}w\right)h_{j},\quad 0<t<1$ (47)
Lemma C.5 (Vector-jacobian product).
Given a linear form $S\in\mathbb{C}^{d\times d}$, the corresponding form $R=E^{\prime}(M)^{\dagger}S$ in $\operatorname{i}\mathfrak{su}(d)$ is given by
$R=\sum_{j=1}^{m}\int_{0}^{1}\operatorname{Re}\operatorname{Tr}(h_{j}v^{\dagger})h_{j}\mathrm{d}t$ (48)
where $v(t)\in\mathbb{C}^{d\times d}$ satisfies
$v(1)=\operatorname{i}U(1)^{\dagger}S,\quad\dot{v}=-2\operatorname{i}\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}(h_{j}v^{\dagger})h_{j}M,\quad 0<t<1$ (49)
We postpone the proofs of these results to Section C.2.2. The calculation relies on the so-called adjoint-state method which we briefly outline and illustrate on a concrete example.
C.2.1 The adjoint state method
Calculating the cost function $J(M)$ involves the implicit solution $U(t)$, to an equation of the form $F(U(\cdot),M)=0$ with parameter $M$. Loosely speaking implicit means that finding the solution $U$ given $M$ is a challenging computational task for which convenient closed form expressions cannot be exploited. In particular finding the gradient of $J$ is challenging because it is not given directly by a formula. The method is often a computationally efficient way to overcome this issue. At the core of the adjoint-state method lies the implicit function theorem and duality.
Method outline.
Suppose that the vector parameter $M$ has $n$ real coordinates and that the cost function is
$J(M)=\ell(x(M))$
where $x(M)\in\mathbb{R}^{N}$ is defined as the unique solution to an implicit equation $F(x(M),g)=0$. Then in response to a perturbation $\delta M$ of the parameter $M$, the first variation of $J$ is
$\delta J=\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}J(M+\varepsilon\delta M)\right|_{\varepsilon=0}=\langle\nabla\ell(x(M)),\delta x\rangle$ (50)
where $\delta x$ is the first variation of $x(M)$ given by
$\delta x=\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}x(M+\varepsilon\delta M)\right|_{\varepsilon=0}$ (51)
We are evaluating a linear form on the implicit solution to a linear equation depending on $\delta M$. To relate $\delta x$ to $\delta M$ we match the first order variations in the equation $F(x(M),M)=0$. This implies that $\delta x$ solves a linear system of equation
$A\delta x+B\delta M=0$ (52)
where $A=\partial_{x}F(x(M),M)$ and $B=\partial_{M}F(x(M),M)$ are the jacobians of $F$ with respect to each arguments. The goal is to rewrite Equation (50) in the form $\delta J=\langle h,\delta M\rangle$ where the vector $h\in\mathbb{R}^{n}$ will be the gradient we seek. This is what enables the introduction of an adjoint state $v\in\mathbb{R}^{N}$ solution to the adjoint linear system
$A^{\dagger}v=\nabla\ell(x(M))$
Because using Equation (52) the first variation of $J$ rewrites
$\delta J=\langle x(M),\delta x\rangle=\langle A^{\dagger}v,\delta x\rangle=\langle v,A\delta x\rangle=\langle v,-B\delta M\rangle$
which means that the gradient of $J$ is the vector
$\boxed{\nabla J(M)=-B^{\dagger}v}$
which depends on $M$ via $B$ and the adjoint state $v$.
Concrete example.
For example put $n=N=2$. Define $\ell(x)=|x|^{2}/2$, the squared euclidean norm of $x$ and
$F(x,M)=\begin{pmatrix}\arctan(x_{1})-M_{1}\\ x_{2}-(x_{1}+M_{2})^{2}\end{pmatrix}$.
If $|M_{1}|<\pi/2$ then the vector $x(M)$ solution to $F(x,M)=0$ is given by
$x(M)=\begin{pmatrix}\tan M_{1}\\ (\tan M_{1}+M_{2})^{2}\end{pmatrix}$
This means that $J(M)=\tan(M_{1})^{2}/2+(\tan(M_{1})+M_{2})^{4}/2$ and so
$\partial_{g_{1}}J(M)$ $=(1+\tan(M_{1})^{2})\left(\tan(M_{1})+2(\tan(M_{1})+M_{2})^{3}\right),$ (53)
$\partial_{g_{2}}J(M)$ $=2(\tan(M_{1})+M_{2})^{3}.$ (54)
Let us obtain the same expression for $\nabla J(M)$ using the adjoint state method described above. With the shorthand notation $y=x(M)$, the coordinates of $y$ are $y_{1}=\tan M_{1}$ and $y_{2}=\tan M_{1}+M_{2}$ and so we have
$\nabla\ell(y)=y,\quad A=\begin{pmatrix}(1+y_{1}^{2})^{-1}&0\\ -2(y_{1}+M_{2})&1\end{pmatrix},\quad B=\begin{pmatrix}-1&0\\ 0&-2(y_{1}+M_{2})\end{pmatrix}$
Therefore the adjoint state $v$ has coordinates $v_{1}=(y_{1}+2y_{1}y_{2}+2M_{2}y_{2})(1+y_{1}^{2})$ and $v_{2}=y_{2}$. Accordingly the gradient $\nabla J(M)$ should be given by
$\nabla J(M)=\begin{pmatrix}v_{1}\\ 2(y_{1}+M_{2})v_{2}\end{pmatrix}$ (55)
and it is straightforward to verify that the vector in Eq. (55) obtained via the adjoint state method is, as expected, equal to the correct gradient whose coordinates are given by Eq. (53).
C.2.2 Proofs of the JVP and VJP expressions
Proof of Lemma C.4 (JVP).
We denote by $\delta U(t)$ the first order variation of the trajectory $U(t)$ solution to (46) in response to a perturbation $\delta M$ of $M$
$\delta U(t)=\left.\frac{\mathrm{d}}{\mathrm{d}\varepsilon}U_{M+\varepsilon\delta M}(t)\right|_{\varepsilon=0}$ (56)
By definition $E^{\prime}(M)\delta M=\delta U(1)$. In Equation (46), keep in mind that $U$ itself depends on $M$. Then differentiating with respect to $M$ the initial condition and both sides of the differential equation yields that the first variation of $U$ is the solution starting at $\delta U(0)=0$ to $\dot{\delta U}=-\operatorname{i}U(A+B)$ where
$A$ $=\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}\left((\delta U)^{\dagger}H_{j}UM+U^{\dagger}H_{j}\delta UM+h_{j}\delta M\right)h_{j},$ (57)
$B$ $=U^{\dagger}H\delta U.$ (58)
We can simplify the $A$ term by observing that $\operatorname{Re}\operatorname{Tr}((\delta U)^{\dagger}H_{j}UM)=\operatorname{Re}\operatorname{Tr}(U^{\dagger}H_{j}\delta UM)$, which means that the two first contributions from the left are equal. Then multiplying $\delta U$ by the inverse propagator $U^{\dagger}$ will cancel the contribution of the $B$ term. Indeed if $w(t)=\operatorname{i}U(t)^{\dagger}\delta U(t)$ then $\dot{w}=\operatorname{i}\dot{U}^{\dagger}\delta U+(A+B)$ and, given the differential equation on $U$, the first term simplifies to $\operatorname{i}\dot{U}^{\dagger}\delta U=-B$. Altogether we see that $\dot{w}=A$. Substituting $\delta U=-\operatorname{i}Uw$ in the $A$ terms yields the Equation (47) on $w$ because $w(0)=0$ since $\delta U(0)=0$. ∎
Proof of Lemma C.5 (VJP).
The aim is to find the Hermitian traceless matrix $R$ such that
$\operatorname{Re}\operatorname{Tr}\left(S^{\dagger}\delta U(1)\right)=\operatorname{Re}\operatorname{Tr}\left(R\delta M\right)$ (59)
for any perturbation $\delta M$ of $M$. Define $w=\operatorname{i}U\delta U$ as in Lemma C.4. Observe that for any function $v$
$\operatorname{Re}\operatorname{Tr}\left(v(1)^{\dagger}w(1)\right)=\int_{0}^{1}\operatorname{Re}\operatorname{Tr}\left(\dot{v}^{\dagger}w+v^{\dagger}\dot{w}\right)\mathrm{d}t$ (60)
by the fundamental theorem of calculus. Using the differential equation on $w$ lets us rewrite Equation (60) as
$\operatorname{Re}\operatorname{Tr}\left(v(1)^{\dagger}w(1)\right)$ $-\int_{0}^{1}\operatorname{Re}\operatorname{Tr}\left((\dot{v}+2\operatorname{i}\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}(v^{\dagger}h_{j})h_{j}M)^{\dagger}w\right)\mathrm{d}t$ (61)
$=$ $\int_{0}^{1}\operatorname{Re}\operatorname{Tr}\left(\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}(v^{\dagger}h_{j})h_{j}\delta M\right)\mathrm{d}t$ (62)
Notice the right-hand side only involves the perturbations $\delta M$ but not $w$. This means that choosing $v$ as the solution to Equation (49), the left-hand side becomes equal to the left-hand side of (59). Then choosing $R$ as in Eq. (48) gives (49) and concludes the proof. ∎
C.3 Optimization routine: natural gradient descent
Consider Problem ($\mathscr{P}_{M}$). Instead of optimizing $|P(M)|$ subject to the constraint $E(M)=U_{f}$, we simplify the problem by optimizing a functional $J(M)=\ell(E(M))$ penalizing the discrepancy to the target $U_{f}$. In other words, we look for a $M$ that is admissible for Problem ($\mathscr{P}_{M}$) but need not be time-optimal. However, this is done using a so-called natural gradient descent (see e.g. [30]), a variant of the Gauss-Newton method for optimization. Its update rule $M\mapsto\hat{M}$ reads
$\hat{M}=M-\rho\left[E^{\prime}\left(M\right)^{\dagger}E^{\prime}\left(M\right)+\alpha I\right]^{-1}\nabla J\left(M\right),\quad\text{where}\quad\nabla J\left(M\right)=E^{\prime}(M)^{\dagger}\nabla\ell(E(M)).$ (63)
In Section C.3.1 below we provide different ways to gain intuition on this update rule. Via the Tikhonov regularization parameter $\alpha$, this method allows for a natural way to penalize the variation of the Frobenius (Hilbert-Schmidt) norm of $M$ and hence of the time $T=|P(M)|$. Put differently, this means that we do not optimize the time directly but rather attempt to find an admissible point $M$ for which $|P(M)|$ is just as large as it needs to be for reaching the target.
We use an optimize-then-discretize approach, also known as indirect method. This means that we first compute a descent direction such as $\nabla J(M)$ and only then introduce a numerical scheme to discretize the equations of motion. This is contrast to the use of an automatic differentiation engine which is part of the discretize-then-optimize framework, also known as direct method.
A proof of concept written in Julia is available on the GitHub repository Magicarp. Alternatively a Python code relying on JAX and the direct method is available on the repository pyMagicarp.
Using a standard gradient descent with line search, infidelity plateaus are observed when the qudit dimension becomes large, say $d\geq 8$. This means that regardless of the number of gradient iterations, the value of $J$ remains above a ”too large” positive value even though the minimal value of $J$ is known to be zero. We resort to natural gradient descent to overcome this issue. This method has proven to be more efficient in practice. This can be understood since it incorporates second order information on the cost functional $J$.
More concretely, the hessian $J^{\prime\prime}(\xi)$ of the composition $J(\xi)=\ell(E(\xi))$ reads $J^{\prime\prime}(\xi)=S+R$ where $S=E^{\prime}(\xi)^{\dagger}\ell^{\prime\prime}(E(\xi))E^{\prime}(\xi)$ and $R$ involves $E^{\prime\prime}(\xi)$ and $\nabla\ell(\xi)$. For convex $\ell$, the matrix $S$ is positive semi-definite while the matrix $R$ need not have a ”sign”. In a way, $R$ is not as well behaved as $S$ for optimization purposes. In a Gauss-Newton method, $R$ is discarded and the matrix $S$ is used. In a (non-regularized) natural gradient method, $R$ is discarded and the matrix $S$ is used as if the hessian of $\ell$ was constant equal to the identity matrix.
C.3.1 Principle of natural gradient descent
Up to choosing an orthonormal basis $(M_{i})_{1\leq i\leq n}$ of $\operatorname{i}\mathfrak{su}(d)$, we may identify $M$ with its coordinates $\xi_{i}=\langle M,M_{i}\rangle$. Accordingly we may view $E$ as a map from $\mathbb{R}^{n}\to\mathbb{C}^{d\times d}$.
Let $J:\mathbb{R}^{n}\to\mathbb{R}$ be the objective function. Assume that it splits into the composition of $E$ with a convex function $\ell:\mathbb{C}^{d\times d}\to\mathbb{R}$, so that $J=\ell\circ E$. In practice $\ell(U)=|U-U_{f}|^{2}$ or the infidelity $\ell(U)=1-|\operatorname{Tr}(UU_{f}^{\dagger})|/d$, although a natural choice the latter is not convex.
Main idea.
The intuition behind splitting $J$ as $\ell$ composed with the model or end-point mapping $E$ is the following: when $\ell$ is convex, optimizing in the $U$ space is in general far easier than optimizing $J$ directly in the $M$ space due to non-convexity. To exploit this idea, the natural gradient step takes into account the model parametrization $M\mapsto U(1)=E(M)$. It projects $\nabla\ell(U(1))$, the ”good” descent direction in $U$ space, onto the range of the linearization of $E$ at $M$. Roughly speaking, the range of $E^{\prime}(M)$ describes the directions of motion locally available around $U(1)=E(M)$ within the state set $S=\{E(M),\,M\in\operatorname{i}\mathfrak{su}(d)\}$. This projection can be seen as preconditioning $\nabla J(M)$, the ”bad” gradient in $M$ space, by the regularized Gram matrix $E^{\prime}(M)^{\dagger}E^{\prime}(M)$ which describes how the parameters or degrees of freedom $M$ offer an ”independent” description of the state set $S$ around the state $E(M)$.
Details.
Fix a penalty $\alpha>0$. Given a point $\xi$ with associated trajectory $U$, we update $\xi$ by a step in the direction $\mu$ with step-size $\rho$
$\delta\xi=\rho\frac{\mu}{|\mu|}.$ (64)
The descent direction $\mu$ is determined as the solution to the Tikhonov-regularized least-squares problem
$\operatorname{\inf}_{\mu\in\mathbb{R}^{n}}|E^{\prime}(\xi)\mu+\nabla\ell(E(\xi))|^{2}+\alpha|\mu|^{2}$ (65)
for which the normal equations are
$\left[E^{\prime}(\xi)^{\dagger}E^{\prime}(\xi)+\alpha I\right]\mu=-\nabla J(\xi).$ (66)
because $\nabla J(\xi)=E^{\prime}(\xi)^{\dagger}\nabla\ell(E(\xi))$. Given the solution $\mu$,the step-size $\rho$ minimizes on $(0,+\infty)$ the map
$\rho\mapsto J\left(\xi+\rho\frac{\mu}{|\mu|}\right)$ (67)
Remark 3.
If $\xi$ is not a critical point of $J$, that is $\nabla J(\xi)\neq 0$ then any step $\mu$ defined by a symmetric semi-positive definite system of linear equation $A\mu=-\nabla J(\xi)$ is a descent direction for $J$. That is for such a $\mu$, $J(\xi+\varepsilon\mu)-J(\xi)=-\varepsilon\langle A\mu,\mu\rangle+o(\varepsilon)$ is negative for small enough $\varepsilon$. In particular this holds for any $\alpha\geq 0$ when $A=E^{\prime}(\xi)^{\dagger}E^{\prime}(\xi)+\alpha I$.
C.3.2 Implementation details.
Algorithm to calculate JVP: $E^{\prime}(M)\delta M$.
Given $M$, we solve forward from $t=0$ to $t=1$ the augmented system $(U,\delta U)$ starting at $(I,0)$ where $\delta U(t)=-\operatorname{i}U(t)w(t)$ is the linearization of the trajectory $U$ in the direction $\delta M$ and $w$ is defined by Equation (47). The value at time $t=1$ of $\delta U$ is the desired quantity $E^{\prime}(M)\delta M$.
Algorithm to calculate VJP: $E^{\prime}(M)^{\dagger}S$.
Given $M$, we solve for the terminal state $U(1)$ and evolve backward from $t=1$ to $t=0$ the augmented system $x=(U,v,z)$ starting at $x_{1}(1)=U(1)$, $x_{2}(1)=\operatorname{i}x_{1}^{\dagger}S$ and $x_{3}(1)=0$. The equation of motion for $x_{1}$ and $x_{2}=v$ are respectively given by (46) and (49) with minus signs to account for backward motion, and that of $x_{3}=z$ is $\dot{z}=-\sum_{j=1}^{m}\operatorname{Re}\operatorname{Tr}(h_{j}v)h_{j}$. In this way the value at time $t=0$ of $z$ is the desired quantity $R=E^{\prime}(M)^{\dagger}S$.
Algorithm to calculate $\nabla J(M)$:
Given $M$, apply (ii) with $S=\nabla\ell(E(M))$.
Algorithm to calculate combined VJP and JVP: $E^{\prime}(M)^{\dagger}E^{\prime}(z)\delta M$.
Find the update step of the natural gradient descent requires solving a symmetric positive-definite linear system involving the gram matrix of $E^{\prime}(M)$ regularized by $\alpha I$. Since $\alpha>0$ and to avoid creating the large non-sparse matrix $E^{\prime}(M)$ of order scaling like $d^{2}$, we appeal to matrix-free methods of Krylov-type. For instance GMRES [34, 38].
Given $\delta M$, we apply (a) to get $S:=E^{\prime}(M)\delta M$ and then (b) to obtain $R=E^{\prime}(M)^{\dagger}S$. Then $R+\alpha\delta M$ is the desired image of $\delta M$ under the linear map defining the normal equations. The step-size $\rho$ is calculated using the golden-section method, a gradient-free optimization method for unimodal functions [33]. This assumption is usually not verified but the method empirically works decently well in our setting.
Initialization.
Whenever $M^{(0)}$ is orthogonal to the linear span of the control Hamiltonians, it is an equilibrium point of the optimization procedure. In particular for some targets $U_{f}$, the reasonable initial guess $M^{(0)}=\operatorname{i}\log(U_{f})$ could be one such equilibrium point.
Accordingly a random initial-guess is preferable without further a priori knowledge. For the purpose of minimizing $T=|P(M)|$ subject to $E(M)=U_{f}$, it is additionally preferable for $|P(M^{(0)})|$ to be relatively ”small” and to choose the penalty $\alpha$ quite large.
Typically $\alpha=10^{-3}$ is a robust choice with respect to the choice of target $U_{f}$ and dimension $d$. If $\alpha$ is larger, we expect $|P(M)|$ to be smaller when the algorithm stops, but this may require a lot more additional iterations.
Validation.
Our stopping criterion is based on the value of the objective function and of the form $J(M)\leq\tau$ where $\epsilon$ is a given absolute tolerance. Once the algorithm stops, we validate the candidate numerical solution $M$ using an independent and high-fidelity ODE solver to compute the $U(1)$ associated to this $M$. We evaluate the corresponding discrepancy to the target $J_{\rm val}(M)$ and check whether it is indeed less then or equal to $\tau$. Typically $\tau=10^{-4}$ if $\ell$ is the infidelity $\ell(x)=1-|\operatorname{Tr}(UU_{f}^{\dagger})|/d$.
(vii) Mesh-refinement.
Given a uniform grid of the interval $[0,1]$ with stepsize $\delta t$, we optimize until the stopping criterion $J(M)\leq\tau$ is met. If the result is not validated in the sense of (f), that is $J_{\rm val}(M)>\tau$, then we refine the mesh. For instance by a factor two $\delta t_{\rm new}=\delta t/2$. Keeping $M$ as initial guess, we then repeat the previous steps until $J_{\rm val}\leq\tau$ or the number of iterations exceeds a pre-defined threshold. This can be seen as a continuation method over the mesh-size.
Appendix D Population dynamics for a QFT on the triple decker
[…]
Figure 12: Controls for the triple decker over time to produce a QFT gate, using the GRD. The controls correspond to addressable energy difference according to links in the graph displayed in fig. 2b. The population’s dynamics is shown in fig. 11. Note that certain controls are unused. Which will be unused depends on the non-unique arbitrary path taken to reach each nodes, which follows a Travelling Salesman algorithm.